OpenStudy (anonymous):

f the supply and demand equations for a product are p=3/100 q+4 and p= 1/100 q+6 respectively where q is the number of units produced and is the price per unit. find a. the equilibium quantity and b. the equilibrium price

OpenStudy (anonymous):

Okay: to find the equilibrium point, you need to find where the two lines intersect. That means that: \[p1 = p2\]\[3q/100 + 4 = q/100 + 6\]\[q/50 = 2\]\[p = 3*100/100 +4 = $7\]\[p = 100/100 +6 = $7\]

OpenStudy (anonymous):

how did you get q/50=2

OpenStudy (anonymous):

subtract q/100 from both sides, you get 2q/100 = 2, or q/50 =2

OpenStudy (anonymous):

ok thanks and both equalibrium is $7?

OpenStudy (anonymous):

price is $7 (I just checked the quantity (q=100)) against both equations

OpenStudy (anonymous):

ok so the quantity is 100